City Street Fund has a portfolio of $432 million and liabilities of $12 million.

a. If there are 28 million shares outstanding, what is net asset value?

Net asset value            $

b-1. If a large investor redeems 2 million shares, what happens to the portfolio value? (Enter your answer in millions.)

Portfolio value (Click to select)increasesdecreases to $  million.

b-2. If a large investor redeems 2 million shares, what happens to shares outstanding? (Enter your answer in millions.)

Shares outstanding (Click to select)increasesdecreases to  million.

b-3. If a large investor redeems 2 million shares, what is net asset value?

Net asset value            $
Answer #1

Net Asset Value = Assets - Liabilities

= 432 million - 12 million

= $420 million

per share = 420/28 = $15

b-1 Portfolio value decreases to 15*26 million = $390 million

b-2Shares outstanding decreases to 26 million

b-3 NAV = 390/26 = $15 per share